Length of a sonometer wire is either 95 cm or 100 cm. In both the cases a tuning fork produces 4 beats then the frequency of tuning fork is?

Grade:12th pass

1 Answers

Nikhil
6 years ago
n=|f1-f2|4=|v/2l-v/2l`|.l=95cm ; l`=100cm.We know that: v= frequency×wavelengthV=15200cm/sHence frequency=152/156Beat frequency